1. Solve the given differential equation by using an appropriate substitution. The DE is a Bernoulli equation

x dy/dx - (1+x)y = x(y)^2


2. Solve the given initial-value problem. The DE is a Bernoulli equation.

x^2 dy/dx - 2xy = 4y^4, y(1)= 1/2
1
Expert's answer
2020-07-20T17:49:13-0400

Question 1: "x\frac{dy}{dx}-(x+1)y=xy^2"

Solution: Dividing equation by "-xy^2" :

"-y^{-2}y'+\frac{x+1}xy^{-1}=-1"

Substituting "z=y^{1-2}=y^{-1}" , "\frac{dz}{dx}=z'=-y^{-2}y'" , therefore: (1)

"z'+\frac{x+1}{x}z=-1"

Substituting "z=uv" , "z'=u'v+uv'" : (2)

"u'v+uv'+\frac{x+1}{x}uv=-1"

"u'v+u(v'+\frac{x+1}{x}v)=-1" (3)

Setting the part inside () to 0 and separating variables:

"v'+\frac{x+1}{x}v=0"

"\frac{dv}{dx}=- \frac{x+1}{x}v"

"\int\frac{dv}v = -\int\frac{x+1}{x}dx"

"ln(v)=-(x+ln(x)+ln(k))"

"v=e^{-(x+ln(x))}=\frac1 {kxe^x}" (4)

Putting "v" from (4) into equation (3):

"\frac1 {kxe^x}u'=-1"

"\int du=-k\int xe^x"

"u=-k(xe^x-e^x+e^x+C)=-k(e^x(x-1)+C)"

Finding z from (2):

"z=uv=\frac{-(e^x(x-1)+C)}{xe^x}"

Finding y from (1):

"y=z^{-1}=-\frac{xe^x}{e^x(x-1)+C}"

Answer 1: "y=-\frac{xe^x}{e^x(x-1)+C}"


Question 2: "x^2y'-2xy=4y^4" , "y(1)=\frac 12"

Solution: Dividing equation by "-\frac1 3 x^2 y^4" :

"-3y^{-4}y'+\frac 6 x y^{-3}=-\frac {12} {x^2}"

Substituting "z=y^{-3}" , "z'=-3y^{-4}y'", therefore: (1)

"z'+\frac6xz=-\frac{12}{x^2}"

Substituting "z=uv , z'=u'v+uv'" : (2)

"u'v+uv'+\frac6xuv=-\frac{12}{x^2}"

"u'v+u(v'+\frac6xv)=-\frac{12}{x^2}" (3)

Setting the part inside () to 0 and separating variables:

"v'+\frac6xv=0" (4)

"v'=-\frac6xv"

"\int \frac {dv} v = -6 \int \frac {dx} x"

"ln(v)= -6 ln(x)-ln(k)=ln(kx^{-6})"

"v=\frac 1 {kx^6}"

Putting "v" from (4) into equation (3):

"\frac 1 {kx^6}u'=-\frac{12}{x^2}"

"\int du = \int -12kx^4dx"

"u= k(-\frac {12x^5}5+C)"

Finding z from (2):

"z=uv=-(\frac {12x^5}5+C)*\frac1{x^6}"

Finding y from (1):

"y=\sqrt[3]{\frac{5x^6}{12x^5+5C}}"

Assuming ​ "y(1)=\frac 12" : "\sqrt[3]{\frac{5*(1)^6}{12*(1)^5+5C}}=\frac12"

"\sqrt[3]{\frac{5}{12+5C}}=\frac12"

"\frac{5}{12+5C}=\frac18"

"C=\frac{52}5"

Substitute C into general solution:

"y=\sqrt[3]{\frac{5x^6}{12x^5+52}}"

Answer 2: "y=\sqrt[3]{\frac{5x^6}{12x^5+52}}"